OpenStudy (anonymous):

simplify the expression -5x+6x+16-11x

OpenStudy (anonymous):

I got -10x+16 ??

jhonyy9 (jhonyy9):

-5+6=1-11=-10 these are coefficients of x so -10x+16 look right sure

OpenStudy (anonymous):

are you saying it " looks right" lol or is right?